# Can Large Language Models Improve Spectral Graph Neural Networks?

### v1

@@ 5.4 Efficiency Studies (RQ3)
To evaluate the efficiency of the proposed method, we measure both its monetary cost and time overhead. Since our method employs low-cost GPT-4o Mini API calls to estimate homophily levels on different datasets, a small monetary expense is incurred. Table 4 reports the input, output, and total monetary cost per dataset. As shown, the cost for each dataset  [^2] Please refer to Appendix C for hyperparameter sensitivity results on additional datasets. remains below \$0.2, indicating that our method achieves high model performance at a minimal financial expense.
To further assess the time efficiency of our approach, we compare the per-epoch and total training time with those of the original polynomial spectral GNNs. As shown in Table 5, the runtime of our method remains similar to the original GNNs across all ten datasets, suggesting that the proposed approach does not increase time complexity.
In practice, the only additional computation introduced by the proposed method, compared to standard spectral GNNs, lies in the construction of heterophily-aware basis functions based on the predicted homophily levels. This step is shared across all graph convolution layers and can thus be precomputed for different polynomial orders. During training, the precomputed basis vectors can be directly loaded to significantly reduce computational overhead. Furthermore, it is worth emphasizing that our approach does not rely on costly fine-tuning or local model deployment, highlighting both the efficiency and practical applicability of the proposed method.
@@ 5.5 Parameter Sensitivity Analysis (RQ4)
In the Methodology Section, we introduce a key hyperparameter $\beta$ to balance the contribution between the original polynomial basis and the heterophily-aware basis constructed using predicted homophily. A lower value of $\beta$ indicates a greater reliance on the heterophily-aware basis, while a higher $\beta$ emphasizes the original polynomial basis. Figure 3 illustrates the sensitivity of hyperparameter $\beta$ on four datasets. As shown, different polynomial models exhibit distinct trends depending on the dataset. For example, BernNetPLUS demonstrates an upward trend in the Washington and Texas datasets. This may be attributed to BernNet’s strong capacity to approximate arbitrary filters, making the original polynomial basis more influential. In contrast, GPRGNNPLUS and JacobiConvPLUS appear largely insensitive to changes in $\beta$, suggesting that both the original and heterophily-aware bases are equally effective at capturing meaningful filters. Conversely, on the Cora and Citeseer datasets, all three models show a generally decreasing trend, indicating that for homophilic graphs, incorporating more heterophily-aware basis functions—constructed based on predicted homophily—can lead to improved performance.
@@ 6 Conclusion
This paper is the first to investigate the potential of LLMs in enhancing spectral GNNs. We propose a novel framework that utilizes LLM-predicted homophily to construct a heterophily basis, which is then integrated into existing spectral GNN architectures. The approach begins by sampling a subset of edges from the graph and employing an LLM to infer whether connected node pairs belong to the same class, thereby estimating homophily. This estimated homophily guides the construction of a heterophily-aware representation, enabling spectral GNNs to handle both homophilic and heterophilic graphs better.
Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of various spectral GNNs, while incurring minimal computational and monetary overhead. These results highlight the promising synergy between LLMs and spectral graph learning, and open up new directions for future research at the intersection of language models and graph neural architectures.

### v2

@@ 5.4 Efficiency Studies (RQ3)
We assess the efficiency of our method in terms of both monetary and time costs. Leveraging low-cost GPT-4o-mini API calls for homophily estimation, the financial overhead remains negligible—under \$0.2 per dataset as shown in Table 4. For time efficiency, Table 5 compares our method with baseline spectral GNNs in per-epoch and total training time, showing no significant runtime increase. The only extra cost stems from constructing heterophily-aware basis functions, which are precomputed once per polynomial order and reused across layers. During training, these basis vectors are simply loaded, adding minimal overhead. Notably, our method requires no fine-tuning or local model deployment, making it lightweight and practical.
@@ 5.5 Parameter Sensitivity Analysis (RQ4)
Our method introduces a key hyperparameter $\beta$ to balance the original polynomial basis and the heterophily-aware basis informed by predicted homophily. Smaller $\beta$ values emphasize the heterophily-aware component, while larger values favor the original basis. Figure 3 shows the performance sensitivity to $\beta$ across four representative datasets. The results reveal diverse trends. For example, in Washington and Texas, BernNetPLUS shows an upward trend, indicating its strong filter capacity benefits from the original basis. In contrast, the other three methods maintain stable performance in Washington and Texas, suggesting robustness to $\beta$ and complementary contributions from both bases. On datasets like Cora and Citeseer, all models perform better with smaller $\beta$, implying that incorporating heterophily-aware basis functions—guided by LLM-estimated homophily—enhances spectral expressiveness even in largely homophilic graphs.
@@ 6 Conclusion
This paper is the first to investigate the potential of LLMs in enhancing spectral GNNs. Specifically, we propose a novel method that utilizes LLM-predicted homophily ratio to construct a heterophily basis, which is then integrated into existing spectral GNN architectures, enabling spectral GNNs to handle both homophilic and heterophilic graphs better. Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of various spectral GNNs, while incurring minimal computational and monetary overhead. Besides, these results highlight the promising synergy between LLMs and spectral graph learning, and open up new directions for future research at the intersection of larger language models and graph neural networks.
Appendix A More Experimental Details
A.1 Datasets
  • Cora, Citeseer, and Pubmed (Guo et al. 2024) are three widely used citation network datasets in graph learning. In each dataset, nodes represent papers and edges indicate citation links between them. Node features are bag-of-words representations of the papers, and labels correspond to document categories.
  • Cornell, Texas, and Wisconsin (Yan et al. 2023) are webpage networks collected from computer science departments of different universities. In these datasets, nodes represent web pages and edges denote hyperlinks between them. We extract node features by encoding the raw webpage content using the pre-trained bert-base-uncased model.
  • Children and History (Yan et al. 2023) are two subsets of the Amazon-Books dataset, containing books labeled with the second-level categories “Children” and “History”, respectively. Nodes represent books, and edges indicate frequent co-purchase or co-view relationships. We use the book titles and descriptions as textual attributes, encoded with the pre-trained roberta-base model.
  • Fitness (Yan et al. 2023) is a subset of the Amazon-Sports dataset, containing items labeled with the second-level category “Fitness”. Nodes represent fitness-related products, and edges indicate frequent co-purchase or co-view relationships. We use the product titles and descriptions as textual features, encoded with the pre-trained roberta-base model.
A.2 Baselines
To facilitate understanding of the models used in our experiments, we briefly introduce the backbone polynomial spectral GNN and the baseline GNN methods evaluated in this paper:
  • GCN (Kipf and Welling 2017) introduces a spectral-based convolution operation on graphs, aggregating information from immediate neighbors with a normalized adjacency matrix.
  • GraphSAGE (Hamilton, Ying, and Leskovec 2017) samples and aggregates features from a node’s local neighborhood using functions like mean or LSTM, enabling inductive learning on large-scale graphs.
  • GAT (Veličković et al. 2019) leverages self-attention mechanisms to assign different importance weights to neighbors, enhancing the flexibility of node representation learning.
  • APPNP (Gasteiger, Bojchevski, and Günnemann 2019) decouples feature transformation and propagation by applying personalized PageRank-based diffusion on top of a neural predictor.
  • GPRGNN (Chien et al. 2021) extends APPNP by directly assigning a learnable coefficient to each order of the normalized adjacency matrix.
  • BernNet (He et al. 2021) expresses the filtering operation with Bernstein polynomials and forces all coefficients to be positive.
  • JacobiConv (Wang and Zhang 2022) proposes a Jacobi basis to adapt a wide range of weight functions due to its orthogonality and flexibility.
  • ChebNetII (He, Wei, and Wen 2022) addresses overfitting and the Runge phenomenon by enhancing the original Chebyshev polynomial approximation, reparameterizing the filter response as a learnable coefficient.
  • TFE-GNN (Duan et al. 2024) introduces a triple filter ensemble to combine low-pass, high-pass, and initial features, enabling adaptive extraction of homophily and heterophily.
  • UniFilter (Huang et al. 2024) unifies heterophilic and homophilic bases to construct a universal polynomial filter that alleviates over-smoothing and over-squashing.
A.3 Hyper-parameter Settings
In all experiments, the number of training epochs is set to 1000, and the learning rate is fixed at 0.0001. A three-layer MLP with a hidden dimension of 256 is used for feature transformation. For each dataset, 100 edges are sampled. The key hyperparameter $\beta$ is tuned over the range $\{0.0,0.1,0.2,\dots,1.0\}$. The optimal $\beta$ values for each method on each dataset are shown in Table 1.
Table 6: Optimal $\beta$ values for each method on each dataset.
Dataset Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
GPRGNNPLUS 0.5 1.0 1.0 1.0 1.0 0.1 0.7 0.6 0.3
BernNetPLUS 0.9 0.5 1.0 0.2 0.1 0.2 0.8 0.3 0.5
JacobiConvPLUS 0.4 0.0 0.0 0.8 0.7 0.6 0.1 0.1 0.1
ChebnetIIPLUS 0.4 0.8 1.0 0.1 0.1 0.0 0.2 0.8 1.0